# Understanding the Role of Training Data in Test-Time Scaling

### v1

@@ Acknowledgments
AJ was supported in part by the Sloan fellowship in mathematics, the NSF Award DMS-2311024, an Amazon Faculty Research Award, an Adobe Faculty Research Award and an iORB grant form USC Marshall School of Business. BM was supported in part by the NSF CAREER Award 2146492, NSF-Simons AI Institute for Cosmic Origins (CosmicAI), NSF AI Institute for Foundations of Machine Learning (IFML), and an Okawa Research Award.
@@ Appendix A Proofs of Theorems and Technical Lemmas
Lemma A.1 is very similar to [8, Lemma C.2]. The difference is that here the features covariance is non-identity, while we do not do CoT during training.
@@
where the penultimate step holds since the cross term is an odd function of $w_{\tau}$ and so its expectation is zero. The other eliminated term is squared and so non-negative. In the last step we used that $w_{\tau}^{\top}\hat{\Lambda}W_{14}$ is scalar and so can be replaced by its transpose. In addition $W_{24}$ is also scalar and can commute with $\hat{\Lambda}$. now observe that $V_{32}$ and $W_{14}$ do not offer any flexibility in minimizing the loss function as their effect can be absorbed in $V_{31}W_{24}$. Therefore, we can set them to zero. Hence,
@@
Define $\tilde{\Delta}_{i}:=\sum_{j=1}^{2k}(\frac{2k\delta}{\sigma_{i}})^{j}$. We next proceed by bounding $\operatorname{\mathbb{E}}(\tilde{\Delta}_{i})$. Observe that $\hat{\Lambda}=\frac{1}{m}\sum_{i\in[m]}x_{i}x_{i}^{\top}$, with $x_{i}\sim{\sf N}(0,\Lambda)$. Using concentration bounds on random matrices with independent sub-gaussian rows (See e.g. [18, Eq. 5.26]), we get that with probability at least $1-2e^{-ct^{2}}$,
@@
where $(b)$ follows from (A.11) and $(c)$ follows from (A.13). In addition, step $(a)$ follows from the following lemma from [7].
@@
([7, Eq.(3.3.39)]) Let $A$ be a given $d$ by $d$ matrix and let $m$ be a given positive integer. For all $p>0$ we have

### v2

@@ Acknowledgments
AJ was supported in part by the Sloan fellowship in mathematics, the NSF Award DMS-2311024, an Amazon Faculty Research Award, an Adobe Faculty Research Award and an iORB grant form USC Marshall School of Business. BM was supported in part by the NSF CAREER Award 2146492, NSF-Simons AI Institute for Cosmic Origins (CosmicAI), NSF AI Institute for Foundations of Machine Learning (IFML).
@@ Appendix A Proofs of Theorems and Technical Lemmas
Lemma A.1 is very similar to (Huang et al., 2025a, Lemma C.2). The difference is that here the features covariance is non-identity, while we do not do CoT during training.
@@
where the penultimate step holds since the cross term is an odd function of $w_{\tau}$ and so its expectation is zero. The other eliminated term is squared and so non-negative. In the last step we used that $w_{\tau}^{\top}\hat{\Lambda}W_{14}$ is scalar and so can be replaced by its transpose. In addition $W_{24}$ is also scalar and can commute with $\hat{\Lambda}$. Note that $V_{32}$ and $W_{14}$ do not offer any flexibility in minimizing the loss function as their effect can be absorbed in $V_{31}W_{24}$. Therefore, we can set them to zero. Hence,
@@
Define $\tilde{\Delta}_{i}:=\sum_{j=1}^{2k}(\frac{2k\delta}{\sigma_{i}})^{j}$. We next proceed by bounding $\operatorname{\mathbb{E}}(\tilde{\Delta}_{i})$. Observe that $\hat{\Lambda}=\frac{1}{m}\sum_{i\in[m]}x_{i}x_{i}^{\top}$, with $x_{i}\sim{\sf N}(0,\Lambda)$. Using concentration bounds on random matrices with independent sub-gaussian rows (See e.g. (Vershynin, 2010, Eq. 5.26)), we get that with probability at least $1-2e^{-ct^{2}}$,
@@
where $(b)$ follows from (A.11) and $(c)$ follows from (A.13). In addition, step $(a)$ follows from the following lemma from Horn and Johnson (1994).
@@
((Horn and Johnson, 1994, Eq.(3.3.39))) Let $A$ be a given $d$ by $d$ matrix and let $m$ be a given positive integer. For all $p>0$ we have
@@ Appendix B Additional Experiments
B.1 Effect of task selection on test time scaling
To demonstrate the improvement we get from our task selection procedure, we consider the set up of Section 5.1, where we generate $w_{\rm test}$ with i.i.d entries from $N(0,1)$. During the test time we initialize with $w_{0}=0$ and let the model generate the final estimate of $w_{\rm test}$ after $k$ step generation. We set the prompt length during training to $n=50$ and prompt length during the test to $m=500$. In the plot below we show how the error $\|w_{\rm test}-w_{k}\|$ behave for the following task selection procedures: 1) Optimal task selection: We set the probabilities $\pi_{\ell}$ by solving the optimization problem (4.5). 2) Uniform selection: We select tasks during training with equal probability. 3) Easy task selection: We select only the easy tasks (Easy-Short or Easy-Long as described in Section 5.1) with equal probabilities. In Figure 5, the estimation error for different task selection procedure is plotted versus $k$, the length of generations during test time before outputting the final estimate.
As we see under the optimal choice the error goes down with $k$, while under under the two other procedures the error goes up with $k$, indicating that a proper choice of tasks during training can avoid overthinking, which can occur under other choices of tasks during training.
Figure 5: Effect of task selection during training on test-time scaling